roots of the equation x2+px+q=0 are 1 and 2. Roots of the equation qx2-px+1=0 must be?
(a)1/2 and 1
(b)1/2 and -1
(c)-1/2 and -1
(d)-1/2 and 1

Grade:12th pass

1 Answers

Khimraj
6 years ago
roots of equation x2+px+q =0 are 1 and 2.
sum of roots of equation = -p = 1+2 =3
So p = -3
product of roots of equation = q = 1*2  =2
so q= 2
Then equation qx2-px +1 =0 becomes 2x2 + 3x+1 =0 
(2x+1)(x+1) = 0
So roots are x = -1/2 and x= -1.
Hope it clears. If you like answer then please approve it.